OpenStudy (anonymous):

sin^2theta+costheta= 2 I have tried using pythagorean identities but I am stuck after removing sin^2 and substituting.

OpenStudy (anonymous):

1-cos^2(theta)+cos(theta)-2=0 cos^2(theta)-cos(theta)+1=0 cos(theta)=\[\frac{ -1\pm \sqrt{1-4} }{ 2 }\] are you sure of the question?

OpenStudy (anonymous):

no, I am positive that this is the question. you have gone more in depth than I think it needs to be. I am supposed to use pythagorean identities. I think. I may be wrong. this is not my strongest area of study.

OpenStudy (anonymous):

maybe I gone too far, but still a contradiction I think, anyways that's my solution I might be wrong :(